# AC-Refiner: Efficient Arithmetic Circuit Optimization Using Conditional Diffusion Models

### v1

@@ I Introduction
The optimization of arithmetic circuits, however, presents significant challenges due to the vast design space, which grows exponentially with input bit width. Over the years, a range of manual design techniques [1, 2, 3, 4, 5] and algorithmic methods [6, 7, 8, 9, 10, 11, 12, 13], which focus on minimizing theoretical properties such as logic depth and size. Despite the efforts, there remains a substantial gap between theoretical optimization and real-world performance. This discrepancy stems from the inherent difficulty of accurately modeling physical implementation constraints (e.g. capacitive loading and routing congestion) at early design stages [14, 15]. Consequently, purely theoretical optimizations cannot guarantee optimal quality-of-results (QoRs) in practice.
@@
On the one hand, reinforcement learning (RL) has been applied to optimize designs through iterative interactions with physical synthesis tools [15, 19, 20, 16]. While RL-based methods demonstrate potential, they often require extensive trial-and-errors before converging on optimized solutions. This iterative process can be prohibitively time-consuming and computationally expensive, particularly for large-scale multipliers [16]. Consequently, RL-agents often struggle to effectively navigate the vast design space within limited budgets, yielding suboptimal design quality.
On the other hand, generative models offer a promising alternatives to address the efficiency limitation of RL. By capturing the structural patterns of prior arithmetic circuit designs, they can rapidly sample diverse promising candidates for parallel evaluation. Prior researches have explored variational autoencoders (VAEs) for prefix adder optimization [17, 18]. However, VAEs exhibit limited capability to model fine-grained structural details [21]. For arithmetic modules with complex design constraints (e.g. multiplier compressor trees), it is difficult for VAEs to sample valid designs.
Figure 1: (a) Image generation with conditional diffusion model. (b) AC-Refiner represents arithmetic circuits similar to images, and conditionally samples circuit designs with desired properties, enabling efficient optimization.
In this paper, we aim to address the limitation of existing learning-based arithmetic circuit optimization, leveraging recent advancement of generative models. To this end, we propose AC-Refiner, which employs conditional diffusion models [22, 23, 24] for the comprehensive and efficient optimization of multipliers, as shown in Fig. 1.
  • Circuit Generation: Unlike VAEs, which generate arithmetic circuit designs in a single inference step, diffusion models employ an iterative inference process. Starting from a random initial design, they progressively refine the structure, ensuring close adherence with design constraints. Incorporated with proper legalization procedures, diffusion models allow AC-Refiner to deliver valid and diverse candidate designs.
  • Guided Optimization: To achieve efficient optimization, AC-Refiner introduces conditional sampling mechanism to explore high-potential designs. Specifically, AC-Refiner adopts neural cost predictors whose gradient steer the inference process toward promising outcomes. Additionally, AC-Refiner leverages discovered high-quality designs to fine-tune the models, which facilitates focused exploration near Pareto frontier.
@@ II-A Multiplier Architecture
    Compressor Tree: As illustrated in Fig. 2(b), the CT compresses PPs for multiple stages until no more than two PPs remain in each column. It is typically composed by various types and quantities of compressors, with full adders (FAs) and half adders (HAs) being predominantly used. These basic compressors take multiple bits (3 for FA and 2 for HA) of the same weight as input, producing one bit of the original weight and another of higher weight.
@@ III-B Model Training
Given a valid CT configuration $\mathcal{T}$, the mapping from $\mathcal{T}$ to RTL implementation is not unique, as input assignments of co-located compressors can be permuted without affecting the multiplication correctness. To obtain a unique and high-quality CT design, we adopt a deterministic interconnection scheme that prioritizes assigning input bits with minimal signal arrival time to FAs over HAs. The internal delay of compressors are estimated on the selected technology node.
Figure 4: Data representation of prefix adder.
@@
    We convert the circuit representation into binary bitmap. For CT designs, we follow Analog-Bit [31] to split integers into multiple binary bits with different weights. The resulting bitmap is denoted as $\mathcal{T}_{B}$, as shown in Fig. 5.
@@
    We cast the input bitmap to a real-valued tensor, denoted as $x_{0}$. Each binary bit $b\in\{0,1\}$ is mapped to a corresponding real value $r\in\{-1.0,1.0\}$. Conversely, a real-valued $r$ can be quantized back to discrete $b$ according to each element’s sign.
@@
Training Cost Predictor. Apart from training diffusion models, AC-Refiner also utilizes neural cost predictors to guide the circuit generation. Specifically, we define function $f$ as the mapping from arithmetic circuit design to the post-synthesis Quality-of-Results (QoR). To enable gradient-based guidance, $f$ takes in the same continuous tensor $x$ as the diffusion model $\epsilon_{\theta}$. To trade off multiple competing QoR metrics and obtain Pareto-optimal circuit design, we conduct multiple synthesis flow to evaluate the arithmetic circuit design $x$. The considered design scenarios include timing-driven, area-driven, and balanced optimization. The total cost $f(x)$ is defined as a weighted sum of delay and area across all the scenarios:
@@
Since $f$ lacks a closed-form expression, we aim to train a differentiable regression model $f_{\pi}$ to approximate the ground truth $f$. We implement $f_{\pi}$ as a 3-layer CNN composed by convolutional residual blocks [33]. Given design $x$ labeled with its corresponding QoR $y$, the training process is formulated as minimizing the following MSE loss:
@@ III-C Conditional Design Generation
Figure 6: Conditional generation of arithmetic circuits. Gradient-guided sampling (Step 1-6) drives the generated designs towards achieving the target quality-of-results (QoR). Self-reflection (Step 7) prevents the gradient guidance from compromising the structrual correctness.
For efficient design optimization, we expect to prune suboptimal candidate designs whose predicted QoR $\tilde{y}=f_{\pi}(\hat{x}_{0})$ does not reach a desirable target $y^{*}$. This objective is formulated as minimizing loss $\mathcal{L}(y^{*},\tilde{y})$, where we select mean-square error as the loss function (Step 3). To achieve effective minimization, we backpropagate the gradient of loss $\mathcal{L}$ down to $x_{t}$ (Steps 4- 5), and add a gradient rectification term $\frac{\partial\mathcal{L}}{\partial x_{t}}$ in the denoising step from $x_{t}$ to $x_{t-1}$ (Step 6):
@@
The resulting intermediate state $x_{t}$ has in-distribution pattern at timestep $t$ and better alignment with the gradient guidance. In plain word, self-reflection mechanism enables selective sampling of $x_{t}$ instances that largely satisfy the gradient guidance. This iterative process is repeated multiple times before advancing to timestep $t-2$.
Algorithm 1 summarizes the overall conditional design generation process, which integrates the proposed gradient-guided sampling and self-reflection mechanism. Notice that Algorithm 1, while largely respecting the design constraints, cannot perfectly guarantee structural integrity, which necessitates the legalization process introduced in Section III-D.
@@ III-D Design Legalization
To ensure the functional correctness of the generated arithmetic circuit designs, AC-Refiner introduces heuristic legalization procedures to eliminate design rule violations (DRVs) in the circuits’ tensor representation. For prefix adders, we adopt the established approach from PrefixRL [15], which iterates through all the prefix nodes and fill the missing lower parents to ensure a valid bitmap $\mathcal{P}$. For compressor trees, GOMIL [12] utilizes integer linear programming (ILP) to produce valid designs satisfying Equation (5)-(9). However, the unpredictable behavior of ILP solver may negate the benefits from conditional circuit generation, rendering it unsuitable in our use case.
Empirically, we observe that CT designs generated by the diffusion models typically exhibit few DRVs, which provides opportunities to correct these errors via efficient heuristics. As shown in Algorithm 2, AC-Refiner adopts CT legalization procedure that takes a series of local adjustments to resolve design rule violations. We denote $E(\mathcal{T})$ as the set of error positions in the compressor tree design $\mathcal{T}$, consisting of over-compression (violating Equation (6)) and under-compression (violating Equation (9)). The algorithm handles over-compression by increasing partial products or reducing compressors (Lines 5-9), and addresses under-compression by adding compressors if needed (Lines 10-14). For quick convergence to a valid configuration, the algorithm greedily selects an action $A^{*}$ from all candidate actions $\mathcal{A}$, which minimizes the size of $E(\mathcal{T})$ (Line 15 to 16). When no action can effectively shrink $E(\mathcal{T})$, we force the legalization process to take a detour, selecting an action that incurs the minimum new errors. In most cases, we observe that Algorithm 2 manages to quickly converge on valid solutions.
@@ III-E Model Fine-Tuning
To further improve optimization efficiency, AC-Refiner fine-tunes the models leveraging explored high-quality designs. This process enables the models to better capture the intricate pattern of optimized designs, which increases the likelihood of discovering even superior ones. As shown in Fig. 3, AC-Refiner performs $M$ rounds of optimization. In each round, the generated designs constitutes a new unlabeled dataset $\mathcal{D}^{\prime}$. A subset of $\mathcal{D}^{\prime}$ is evaluated for QoR with physical synthesis tools, creating a labeled dataset $\hat{\mathcal{D}}^{\prime}$. The newly acquired datasets are then merged with the initial dataset $\mathcal{D}$ and $\hat{\mathcal{D}}$ to fine-tune the diffusion model $\epsilon_{\theta}$ and regression model $f_{\pi}$.
@@ IV-A Setup
Platform. All experiments are conducted on a Linux-based platform equipped with a 2.9GHz 64-core Intel Xeon Gold 6226R CPU, 376GB of memory, and an NVIDIA A6000 GPU. We use open-sourced logic synthesis tool Yosys (verision 0.27) [35] and the physical synthesis tool OpenROAD (version 2.0) [36] to assess the Quality-of-Results (QoR) of multiplier designs using the NanGate 45nm technology process library [37].
@@
Hyperparameter Settings. We implement the diffusion models $\epsilon_{\theta}$ and cost predictors $f_{\pi}$ with PyTorch. To evaluate the QoR cost function as in Equation (11), we synthesize multiplier designs under area-driven scenario and timing-driven scenario, and normalize the delay and area with corresponding values of Wallace multiplier [4]. The trade-off weight is set to $w=0.66$. For gradient-guided sampling, we set the target QoR objective $y=0.7$, the guidance strength $s(t)=10\sqrt{1-\alpha_{t}}$, and the self-reflection step to $k=25$. We conduct comprehensive optimization of multipliers covering both CT and CPA. We commence by optimizing CT and fix CPA with the default adder from the synthesis tool. Then we adopt the best explored CT and refine the CPA. For CT and CPA respectively, the initial dataset consists of $|D_{0}|\approx 15000$ unlabeled samples, and $|\hat{D}_{0}|=1000$ designs labeled with post-synthesis QoR. Optimization runs for $M=5$ rounds, where in each round we sample $|D_{m}|=1000$ designs and randomly select $|\hat{D}_{m}|=100$ designs for physical synthesis. In total, the synthesis tools are invoked 3000 times, which is the same as learning-based baselines RL-MUL and ArithmeticTree.
@@ IV-B Multiplier Optimization Result
We compare the Pareto frontiers obtained using different optimization algorithms. Following established evaluation methodologies [19, 16], we select the optimal multiplier from each approach and synthesize them with target delay ranging from 0ns to 2ns to demonstrate the trade-off between timing and area. As shown in Fig. 7, AC-Refiner achieves superior Pareto optimality over all baselines. Both manually designed Wallace multiplier and algorithmic method GOMIL fail to account for the complexities inherent in physical implementation, leading to suboptimal post-synthesis delay and area. RL-MUL modifies only the column-wise compressor count without considering detailed stage-specific compressor assignment or CPA optimization, yielding inferior performance. ArithmeticTree trains the RL agents to construct multipliers in a gate-by-gate manner. Despite fine-grained optimization, it suffers from exploration inefficiency as the RL agents must rely on extensive trial-and-error iterations. Compared with the state-of-the-art ArithmeticTree method, AC-Refiner achieves a delay reduction of up to 15% and an area reduction of up to 10%.
We also examine the QoR improvement by progressively incorporating diffusion-driven optimization. Fig. 8 shows that CTs and CPAs optimized by AC-Refiner (Opt) achieves better QoR over the ones in the initial dataset (Init). Notably, the initial CPAs may not integrate well with the optimized CT than the default adders from synthesis tool (Default), yet AC-Refiner still identifies CPAs that enhance the overall QoR of multipliers.
IV-C Ablations
@@
TABLE I: Systolic array comparison on delay, area, and ADP.
@@
Delay (ns) Area ($\mu\text{m}^{2}$) ADP ($\mu\text{m}^{2}$·ns) Delay (ns) Area ($\mu\text{m}^{2}$) ADP ($\mu\text{m}^{2}$·ns)
@@
Gradient-Guided Sampling. The QoR outcome $\tilde{y}$ of conditional sampling process is influenced by two factors: target QoR $y^{*}$ and guidance strength $s(t)$. We demonstrate their individual effects through the optimization of compressor trees across various bit-width.
In Fig. 9 we examine the impact of $y^{*}$ on $\tilde{y}$. When $y^{*}$ falls in the range of initial dataset, the real QoR $\tilde{y}$ shows strong correlation with $y^{*}$. This phenomenon shows that the neural cost predictor successfully captures the nuanced interplay between circuit structure and physical synthesis, offering effective guidance during conditional circuit generation. When $y^{*}$ is set extremely low, the real QoR outcome generalizes beyond the minima of initial dataset, highlighting the efficacy of diffusion-driven optimization in expanding Pareto frontiers.

### v2

@@ I Introduction
The optimization of arithmetic circuits, however, presents significant challenges due to the vast design space, which grows exponentially with input bit width. Over the years, a range of manual design techniques [1, 2, 3, 4, 5] and algorithmic methods [6, 7, 8, 9, 10, 11, 12, 13], which focus on minimizing theoretical properties such as logic depth and size. Despite these efforts, a substantial gap remains between theoretical optimization and real-world performance. This discrepancy stems from the inherent difficulty of accurately modeling physical implementation constraints (e.g. capacitive loading and routing congestion) at early design stages [14, 15]. Consequently, purely theoretical optimizations cannot guarantee optimal quality-of-results (QoRs) in practice.
@@
On the one hand, reinforcement learning (RL) has been applied to optimize designs through iterative interactions with physical synthesis tools [15, 19, 20, 16]. While RL-based methods demonstrate potential, they often require extensive trial-and-error before converging on optimized solutions. This iterative process can be prohibitively time-consuming and computationally expensive, particularly for large-scale multipliers [16]. Consequently, RL-agents often struggle to effectively navigate the vast design space within limited budgets, yielding suboptimal design quality.
On the other hand, generative models offer a promising alternative to address the efficiency limitation of RL. By capturing the structural patterns of prior arithmetic circuit designs, they can rapidly sample diverse promising candidates for parallel evaluation. Prior research has explored variational autoencoders (VAEs) for prefix adder optimization [17, 18]. However, VAEs exhibit limited capability to model fine-grained structural details [21]. For arithmetic modules with complex design constraints (e.g. multiplier compressor trees), it is difficult for VAEs to sample valid designs.
Figure 1: (a) Image generation with a conditional diffusion model. (b) AC-Refiner represents arithmetic circuits similar to images, and conditionally samples circuit designs with desired properties, enabling efficient optimization.
In this paper, we aim to address the limitations of existing learning-based arithmetic circuit optimization by leveraging recent advancements of generative models. To this end, we propose AC-Refiner, which employs conditional diffusion models [22, 23, 24] for the comprehensive and efficient optimization of multipliers, as shown in Fig. 1. The key contributions of AC-Refiner are summarized as follows:
  • Circuit Generation: Unlike VAEs, which generate arithmetic circuit designs in a single inference step, diffusion models employ an iterative inference process. Starting from a random initial design, they progressively refine the structure, ensuring close adherence to design constraints. Incorporated with proper legalization procedures, diffusion models allow AC-Refiner to deliver valid and diverse candidate designs.
  • Guided Optimization: To achieve efficient optimization, AC-Refiner introduces conditional sampling mechanism to explore high-potential designs. Specifically, AC-Refiner adopts neural cost predictors whose gradient steer the inference process toward promising outcomes. Additionally, AC-Refiner leverages discovered high-quality designs to fine-tune the models, which facilitates focused exploration near the Pareto frontier.
@@ II-A Multiplier Architecture
    Compressor Tree: As illustrated in Fig. 2(b), the CT compresses PPs for multiple stages until no more than two PPs remain in each column. It is typically composed of various types and quantities of compressors, with full adders (FAs) and half adders (HAs) being predominantly used. These basic compressors take multiple bits (3 for FA and 2 for HA) of the same weight as input, producing one bit of the original weight and another of higher weight.
@@ III-B Model Training
Given a valid CT configuration $\mathcal{T}$, the mapping from $\mathcal{T}$ to RTL implementation is not unique, as input assignments of co-located compressors can be permuted without affecting the multiplication correctness. To obtain a unique and high-quality CT design, we adopt a deterministic interconnection scheme that prioritizes assigning input bits with minimal signal arrival time to FAs over HAs. The internal delay of compressors is estimated on the selected technology node.
@@
    We convert the circuit representations into binary bitmaps. For CT designs, we follow Analog-Bit [31] to split integers into multiple binary bits with different weights. The resulting bitmap is denoted as $\mathcal{T}_{B}$, as shown in Fig. 5.
@@
    We cast the input bitmap to a real-valued tensor, denoted as $x_{0}$. Each binary bit $b\in\{0,1\}$ is mapped to a corresponding real value $r\in\{-1.0,1.0\}$. Conversely, a real-valued $r$ can be quantized back to a discrete $b$ according to each element’s sign.
@@
Training Cost Predictor. Apart from training diffusion models, AC-Refiner also utilizes neural cost predictors to guide the circuit generation. Specifically, we define function $f$ as the mapping from an arithmetic circuit design to the post-synthesis Quality-of-Results (QoR). To enable gradient-based guidance, $f$ takes in the same continuous tensor $x$ as the diffusion model $\epsilon_{\theta}$. To trade off multiple competing QoR metrics and obtain Pareto-optimal circuit design, we conduct multiple synthesis flows to evaluate the arithmetic circuit design $x$. The considered design scenarios include timing-driven, area-driven, and balanced optimization. The total cost $f(x)$ is defined as a weighted sum of delay and area across all the scenarios:
@@
Since $f$ lacks a closed-form expression, we aim to train a differentiable regression model $f_{\pi}$ to approximate the ground truth $f$. We implement $f_{\pi}$ as a 3-layer CNN composed of convolutional residual blocks [33]. Given design $x$ labeled with its corresponding QoR $y$, the training process is formulated as minimizing the MSE loss:
@@ III-C Conditional Design Generation
Figure 6: Conditional generation of arithmetic circuits. Gradient-guided sampling (Step 1-6) drives the generated designs towards achieving the target quality-of-results (QoR). Self-reflection (Step 7) prevents the gradient guidance from compromising the structural correctness.
[…]
@@
The resulting intermediate state $x_{t}$ has an in-distribution pattern at timestep $t$ and better alignment with the gradient guidance. In plain words, self-reflection mechanism enables selective sampling of $x_{t}$ instances that largely satisfy the gradient guidance. This iterative process is repeated multiple times before advancing to timestep t2t-2.
Algorithm 1 summarizes the overall conditional design generation process, which integrates the proposed gradient-guided sampling and self-reflection mechanism. While it largely respects the design constraints, it cannot perfectly guarantee structural integrity, which necessitates the legalization process introduced in Section III-D.
@@ III-D Design Legalization
To ensure the functional correctness of the generated arithmetic circuit designs, AC-Refiner introduces heuristic legalization procedures to eliminate design rule violations (DRVs) in the circuits’ tensor representation. For prefix adders, we adopt the established approach from PrefixRL [15], which iterates through all the prefix nodes and fills the missing lower parents to ensure a valid bitmap $\mathcal{P}$. For compressor trees, GOMIL [12] utilizes integer linear programming (ILP) to produce valid designs satisfying Equation (5)-(9). However, the unpredictable behavior of ILP solver may negate the benefits from conditional circuit generation, rendering it unsuitable in our use case.
Empirically, we observe that CT designs generated by the diffusion models typically exhibit few DRVs, which provide opportunities to correct these errors via efficient heuristics. As shown in Algorithm 2, AC-Refiner adopts CT legalization procedure that takes a series of local adjustments to resolve design rule violations. We denote $E(\mathcal{T})$ as the set of error positions in the compressor tree design $\mathcal{T}$, consisting of over-compression (violating Equation (6)) and under-compression (violating Equation (9)). The algorithm handles over-compression by increasing partial products or reducing compressors (Lines 5-9), and addresses under-compression by adding compressors if needed (Lines 10-14). For quick convergence to a valid configuration, the algorithm greedily selects an action $A^{*}$ from all candidate actions $\mathcal{A}$, which minimizes the size of $E(\mathcal{T})$ (Line 15 to 16). When no action can effectively shrink $E(\mathcal{T})$, we force the legalization process to take a detour, selecting an action that incurs the minimum new errors. In most cases, we observe that Algorithm 2 manages to quickly converge on valid solutions.
@@ III-E Model Fine-Tuning
To further improve optimization efficiency, AC-Refiner fine-tunes the models leveraging the explored high-quality designs. This process enables the models to better capture the intricate pattern of optimized designs, which increases the likelihood of discovering even superior ones. As shown in Fig. 3, AC-Refiner performs $M$ rounds of optimization. In each round, the generated designs constitute a new unlabeled dataset $\mathcal{D}^{\prime}$. A subset of $\mathcal{D}^{\prime}$ is evaluated for QoR with physical synthesis tools, creating a labeled dataset $\hat{\mathcal{D}}^{\prime}$. The newly acquired datasets are then merged with the initial dataset $\mathcal{D}$ and $\hat{\mathcal{D}}$ to fine-tune the diffusion model $\epsilon_{\theta}$ and regression model $f_{\pi}$.
@@ IV-A Setup
Platform. All experiments are conducted on a Linux-based platform equipped with a 2.9GHz 64-core Intel Xeon Gold 6226R CPU, 376GB of memory, and an NVIDIA A6000 GPU. We use open-sourced logic synthesis tool Yosys (version 0.27) [35] and the physical synthesis tool OpenROAD (version 2.0) [36] to assess the Quality-of-Results (QoR) of multiplier designs using the NanGate 45nm technology process library [37].
@@
Hyperparameter Settings. We implement the diffusion models $\epsilon_{\theta}$ and cost predictors $f_{\pi}$ with PyTorch. To evaluate the QoR cost function as in Equation (11), we synthesize multiplier designs under area-driven scenario and timing-driven scenario, and normalize the delay and area with corresponding values of Wallace multiplier [4]. The trade-off weight is set to $w=0.66$. For gradient-guided sampling, we set the target QoR objective $y=0.7$, the guidance strength $s(t)=10\sqrt{1-\alpha_{t}}$, and the self-reflection step to $k=25$. We conduct a comprehensive optimization of multipliers covering both CT and CPA. We commence by optimizing CT and fix CPA with the default adder from the synthesis tool. Then we adopt the best explored CT and refine the CPA. For CT and CPA, respectively, the initial dataset consists of $|D_{0}|\approx 15000$ unlabeled samples, and $|\hat{D}_{0}|=1000$ designs labeled with post-synthesis QoR. Optimization runs for $M=5$ rounds, where in each round we sample $|D_{m}|=1000$ designs and randomly select $|\hat{D}_{m}|=100$ designs for physical synthesis. In total, the synthesis tools are invoked 3000 times, which is the same as learning-based baselines RL-MUL and ArithmeticTree.
@@ IV-B Multiplier Optimization Result
We compare the Pareto frontiers obtained using different optimization algorithms. Following established evaluation methodologies [19, 16], we select the optimal multiplier from each approach and synthesize them with a target delay ranging from 0ns to 2ns to demonstrate the trade-off between timing and area. As shown in Fig. 7, AC-Refiner achieves superior Pareto optimality over all baselines. Both the manually designed Wallace multiplier and the algorithmic method GOMIL fail to account for the complexities inherent in physical implementation, leading to suboptimal post-synthesis delay and area. RL-MUL modifies only the column-wise compressor count without considering detailed stage-specific compressor assignment or CPA optimization, yielding inferior performance. ArithmeticTree trains the RL agents to construct multipliers in a gate-by-gate manner. Despite fine-grained optimization, it suffers from exploration inefficiency as the RL agents must rely on extensive trial-and-error iterations. Compared with the state-of-the-art ArithmeticTree method, AC-Refiner achieves a delay reduction of up to 15% and an area reduction of up to 10%.
@@
TABLE I: Systolic array comparison on delay, area, and area-delay product (ADP).
@@
Delay (ns) Area ($\mu\text{m}^{2}$) ADP ($\mu\text{m}^{2}\cdotns$) Delay (ns) Area ($\mu\text{m}^{2}$) ADP ($\mu\text{m}^{2}\cdotns$)
@@
We also examine the QoR improvement by progressively incorporating diffusion-driven optimization. Fig. 8 shows that CTs and CPAs optimized by AC-Refiner (Opt) achieve better QoR over the ones in the initial dataset (Init). Notably, the initial CPAs may not integrate well with the optimized CT than the default adders from the synthesis tool (Default), yet AC-Refiner still identifies CPAs that enhance the overall QoR of multipliers.
IV-C Ablations
Gradient-Guided Sampling. The QoR outcome $\tilde{y}$ of the conditional sampling process is influenced by two factors: target QoR $y^{*}$ and guidance strength $s(t)$. We demonstrate their individual effects through the optimization of compressor trees across various bit-width.
In Fig. 9 we examine the impact of $y^{*}$ on $\tilde{y}$. When $y^{*}$ falls in the range of the initial dataset, the real QoR $\tilde{y}$ shows a strong correlation with $y^{*}$. This phenomenon shows that the neural cost predictor successfully captures the nuanced interplay between circuit structure and physical synthesis, offering effective guidance for conditional circuit generation. When $y^{*}$ is set extremely low, the real QoR outcome generalizes beyond the minima of the initial dataset, highlighting the efficacy of diffusion-driven optimization in expanding Pareto frontiers.